OpenStudy (anonymous):

A person purchased 5k + 2 items for a total cost of 35k^2 + 29k + 6. Find the average cost per item of this purchase.

OpenStudy (nurali):

\[C _{ave}=\frac{ 35k^2 + 29k + 6 }{ 5k + 2 }\] \[C _{ave}=\frac{(7k+3)( 5k + 2) }{ (5k + 2) }\] \[C _{ave}=7k+3\]

OpenStudy (nurali):

7k+3 is the average cost per item of this purchase.

OpenStudy (nurali):

if correct please click best response.